# Improving LLM Unlearning Robustness via Random Perturbations

### v3

@@ Appendix B Proofs
Consider the output representation of the predicted token $x^{r}_{i}$ given the perturbed retain-query prefix $x^{r,\text{per}}_{<i}$ in the unlearned model $f^{u}(x^{r}_{i}|x^{r,\text{per}}_{<i})$. We show the claim by using the framework of the generative latent variable model (GLVM). Specifically, model $f^{u}$ generates token $x^{r}_{i}$ conditioned on a latent variable $\mathbf{z}^{r,\text{per}}_{<i}$ corresponding to the perturbed prefix $x^{r,\text{per}}_{<i}$, denoted as $f^{u}(x^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i})$. Under Assumption 4.1, the following holds:
$f^{u}(x^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i})=f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon})$ (16)
Since $\bm{\epsilon}$ is small, we approximate the function $f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon})$ around $\mathbf{z}^{r}_{<i}$ by using the first-order Taylor approximation:
$f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon})$ $\approx f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i})+\nabla_{\mathbf{z}^{r}_{<i}}f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i})^{\top}\bm{\epsilon}$ (17)
$f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon})-f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i})$ $\approx\nabla_{\mathbf{z}^{r}_{<i}}f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i})^{\top}\bm{\epsilon}$ (18)
Let $\Delta=f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon})-f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i})$, given that $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\eta\bm{I})$, by the affine transformation of Gaussian variables, we obtain $\Delta\sim\mathcal{N}(\bm{0},\eta\bm{J}^{\top}\bm{J})$, where $\bm{J}=\nabla_{\mathbf{z}^{r}_{<i}}f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i})$ is the Jacobian of $f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i})$ with respect to $\mathbf{z}^{r}_{<i}$. ∎
@@
Let us consider the generation of $x^{r}_{i}$ through the lens of a GLVM. The loss of $x^{r}_{i}$ given the latent representation $\mathbf{z}^{r,\text{per}}_{<i}$ of the prefix $x^{r,\text{per}}_{<i}$ in unlearned model $f^{u}$, is denoted by $f^{u}(x^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i})$. Under Assumption 4.1, the following holds:
$\mathcal{J}(f^{u}(x^{r}_{i}$ $|\mathbf{z}^{r,\text{per}}_{<i}))=\mathcal{J}(f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon}))$ (19)
Since $\bm{\epsilon}$ is small, we linearly approximate function $\mathcal{J}(f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon})$ around $\mathbf{z}^{r}_{<i}$ by using the first-order Taylor approximation:
$\mathcal{J}(f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon}))$ $\approx\mathcal{J}(f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i}))+\nabla_{\mathbf{z}^{r}_{<i}}\mathcal{J}(f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i}))^{\top}\bm{\epsilon}$ (20)
@@
$\Delta\mathcal{J}^{u}\approx\nabla_{\mathbf{z}^{r}_{<i}}\mathcal{J}(f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i}))^{\top}\bm{\epsilon}$ (21)
Under Assumption 4.1 and Assumption 5.1, $\mathcal{J}(f^{\text{rna}}(x^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i}))$ and $\mathcal{J}(f^{\text{rna}}(x^{r}_{i}|\mathbf{z}^{r}_{<i}))$ can be expressed as:
$\mathcal{J}(f^{\text{rna}}(x^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i}))$ $=\mathcal{J}(f^{\text{rna}}(x^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon}))$ (22)
$\approx\mathcal{J}(f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon}+\bm{\delta}_{1}))$ (23)
$\approx\mathcal{J}(f^{u}(x^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i}+\bm{\delta}_{1}))$ (24)
$\approx\mathcal{J}(f^{u}(x^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i}))+\nabla_{\mathbf{z}^{r,\text{per}}_{<i}}\mathcal{J}(f^{u}(x^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i}))^{\top}\bm{\delta}_{1}$ (25)
$\mathcal{J}(f^{\text{rna}}(x^{r}_{i}|\mathbf{z}^{r}_{<i}))$ $\approx\mathcal{J}(f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i})+\bm{\delta}_{2}))$
$\approx\mathcal{J}(f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i}))+\nabla_{\mathbf{z}^{r}_{<i}}\mathcal{J}(f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i}))^{\top}\bm{\delta}_{2}$ (26)
Substituting Eqn. 25 and Eqn. 26, the change in loss in RNA model $f^{\text{rna}}$ of predicted token $x^{r}_{i}$ is approximately:
@@
where $\bm{g}^{\text{per}}=\nabla_{\mathbf{z}^{r,\text{per}}_{<i}}\mathcal{J}(f^{u}(x^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i}))$ and $\bm{g}=\nabla_{\mathbf{z}^{r}_{<i}}\mathcal{J}(f^{u}(x^{r}_{i}|\mathbf{z}^{r}_{<i}))$.
The ratio of the RNA loss change to the original unlearned model loss change is:
@@
$(\bm{g}^{\text{per}})^{\top}\bm{\delta}_{1}-\bm{g}^{\top}\bm{\delta}_{2}$ $\sim\mathcal{N}(0,\nu(||\bm{g}^{\text{per}}||_{2}^{2}+||\bm{g}||_{2}^{2}))$
$\bm{g}^{\top}\bm{\epsilon}$ $\sim\mathcal{N}(0,\eta||\bm{g}||_{2}^{2})$
@@
The ratio of two random normally distributed variables $\frac{(\bm{g}^{\text{per}})^{\top}\bm{\delta}_{1}-\bm{g}^{\top}\bm{\delta}_{2}}{\bm{g}^{\top}\bm{\epsilon}}$ follows a Cauchy distribution with location parameter $x_{0}=0$ and scale parameter $\gamma=\sqrt{\frac{\nu}{\eta}}\left(1+\frac{||\bm{g}^{\text{per}}||_{2}}{||\bm{g}||_{2}}\right)$. The cumulative distribution function of $\text{Cauchy}\left(0,\sqrt{\frac{\nu}{\eta}}\left(1+\frac{||\bm{g}^{\text{per}}||_{2}}{||\bm{g}||_{2}}\right)\right)$ given by
$F(x;x_{0},\gamma)=\frac{1}{2}+\frac{1}{\pi}\arctan\left(\frac{x}{\sqrt{\frac{\nu}{\eta}}\left(1+\frac{||\bm{g}^{\text{per}}||_{2}}{||\bm{g}||_{2}}\right)}\right)$
@@
$\mathbb{P}\left[\frac{(\bm{g}^{\text{per}})^{\top}\bm{\delta}_{1}-\bm{g}^{\top}\bm{\delta}_{2}}{\bm{g}^{\top}\bm{\epsilon}}\leq-1\right]$ $=F(x=-1;x_{0},\gamma)$ (30)
$=\frac{1}{2}+\frac{1}{\pi}\arctan\left(\frac{-1}{\sqrt{\frac{\nu}{\eta}}\left(1+\frac{||\bm{g}^{\text{per}}||_{2}}{||\bm{g}||_{2}}\right)}\right)$ (31)
$=\frac{1}{2}-\frac{1}{\pi}\arctan\left[\sqrt{\frac{\eta}{\nu}}\left(1+\frac{||\bm{g}^{\text{per}}||_{2}}{||\bm{g}||_{2}}\right)^{-1}\right]$ (32)
Appendix C Empirical Validation of Section 3
In this section, we aim to show that the PO forgetting process (minimizing the forget-loss) can be interpreted as injecting random noise into the latent representations of forget-samples during fine-tuning.
@@ Appendix C Empirical Validation of Section 3
We formalize the forgetting through the lens of noise sensitivity [1]. Let $\mathbf{z}^{f}\in\mathbb{R}^{d_{l}}$ be the hidden states vector of forget-sample $x^{f}$ at layer $l$ in the model $f$, where $d_{l}$ is the dimension of layer $l$. Let $g$ be the $(l+1)$-th transformer layer in model $f^{u}$. Consider a random perturbation $\bm{\xi}$ drawn from a Normal distribution $\mathcal{N}(\bm{0},\bm{I})$. The noise sensitivity of $g$ with respect to $\mathcal{N}(\bm{0},\bm{I})$ on forget-set $\mathcal{D}_{f}$, is defined as:
@@
For all unlearned models, we perform grid search for $g$ from the first to the last layer in the model. We use the WMDP-Biology forget-set to compute the noise sensitivity of layers by Eqn. 33. The max sequence length of each forget-sample is set to $512$.
@@ Results.
Discussion.
We employ the noise sensitivity to validate the analysis in Section 3. However, we believe that this definition has broader potential applications. One could explore the noise sensitivity as a metric for measuring unlearning difficulty. This definition generalizes two perspectives: model difficulty and data difficulty. From the model perspective, noise sensitivity can help characterize the unlearning difficulty of specific components—such as an intermediate layer (as described in Eqn. 33), a group of layers, an entire model (e.g., Llama vs. Mistral), or more fine-grained modules in the layer such as MLP, attention patterns, or individual neurons. From the data perspective, the noise sensitivity can be used to evaluate unlearning difficulty at the level of individual samples, sub-classes, or data subsets. We leave these promising directions for future work.

### v4

@@ Appendix B Proofs
Consider the output representation of the predicted token $\mathbf{x}^{r}_{i}$ given the perturbed retain-query prefix $\mathbf{x}^{r,\text{per}}_{<i}$ in the unlearned model $f^{u}(\mathbf{x}^{r}_{i}|\mathbf{x}^{r,\text{per}}_{<i})$. We show the claim by using the framework of the generative latent variable model (GLVM). Specifically, model $f^{u}$ generates token $\mathbf{x}^{r}_{i}$ conditioned on a latent variable $\mathbf{z}^{r,\text{per}}_{<i}$ corresponding to the perturbed prefix $\mathbf{x}^{r,\text{per}}_{<i}$, denoted as $f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i})$. Under Assumption 4.1, the following holds:
$f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i})=f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon})$ (16)
Since $\bm{\epsilon}$ is small, we approximate the function $f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon})$ around $\mathbf{z}^{r}_{<i}$ by using the first-order Taylor approximation:
$f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon})$ $\approx f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i})+\nabla_{\mathbf{z}^{r}_{<i}}f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i})^{\top}\bm{\epsilon}$ (17)
$f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon})-f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i})$ $\approx\nabla_{\mathbf{z}^{r}_{<i}}f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i})^{\top}\bm{\epsilon}$ (18)
Let $\Delta=f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon})-f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i})$, given that $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\eta\bm{I})$, by the affine transformation of Gaussian variables, we obtain $\Delta\sim\mathcal{N}(\bm{0},\eta\bm{J}^{\top}\bm{J})$, where $\bm{J}=\nabla_{\mathbf{z}^{r}_{<i}}f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i})$ is the Jacobian of $f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i})$ with respect to $\mathbf{z}^{r}_{<i}$. ∎
@@
Let us consider the generation of $\mathbf{x}^{r}_{i}$ through the lens of a GLVM. The loss of $\mathbf{x}^{r}_{i}$ given the latent representation $\mathbf{z}^{r,\text{per}}_{<i}$ of the prefix $\mathbf{x}^{r,\text{per}}_{<i}$ in unlearned model $f^{u}$, is denoted by $f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i})$. Under Assumption 4.1, the following holds:
$\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}$ $|\mathbf{z}^{r,\text{per}}_{<i}))=\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon}))$ (19)
Since $\bm{\epsilon}$ is small, we linearly approximate function $\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon})$ around $\mathbf{z}^{r}_{<i}$ by using the first-order Taylor approximation:
$\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon}))$ $\approx\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}))+\nabla_{\mathbf{z}^{r}_{<i}}\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}))^{\top}\bm{\epsilon}$ (20)
@@
$\Delta\mathcal{J}^{u}\approx\nabla_{\mathbf{z}^{r}_{<i}}\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}))^{\top}\bm{\epsilon}$ (21)
Under Assumption 4.1 and Assumption 5.1, $\mathcal{J}(f^{\text{rna}}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i}))$ and $\mathcal{J}(f^{\text{rna}}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}))$ can be expressed as:
$\mathcal{J}(f^{\text{rna}}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i}))$ $=\mathcal{J}(f^{\text{rna}}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon}))$ (22)
$\approx\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}+\bm{\epsilon}+\bm{\delta}_{1}))$ (23)
$\approx\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i}+\bm{\delta}_{1}))$ (24)
$\approx\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i}))+\nabla_{\mathbf{z}^{r,\text{per}}_{<i}}\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i}))^{\top}\bm{\delta}_{1}$ (25)
$\mathcal{J}(f^{\text{rna}}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}))$ $\approx\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i})+\bm{\delta}_{2}))$ (26)
$\approx\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}))+\nabla_{\mathbf{z}^{r}_{<i}}\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}))^{\top}\bm{\delta}_{2}$ (27)
Substituting Eqn. 25 and Eqn. 27, the change in loss in RNA model $f^{\text{rna}}$ of predicted token $\mathbf{x}^{r}_{i}$ is approximately:
$\mathcal{J}(f^{\text{rna}}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i}))-\mathcal{J}(f^{\text{rna}}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}))$ $\approx\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i}))-\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}))$
$+\nabla_{\mathbf{z}^{r,\text{per}}_{<i}}\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i}))^{\top}\bm{\delta}_{1}-\nabla_{\mathbf{z}^{r}_{<i}}\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}))^{\top}\bm{\delta}_{2}$ (28)
@@
where $\bm{g}^{\text{per}}=\nabla_{\mathbf{z}^{r,\text{per}}_{<i}}\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r,\text{per}}_{<i}))$ and $\bm{g}=\nabla_{\mathbf{z}^{r}_{<i}}\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{z}^{r}_{<i}))$.
From Eqn. 21 and Eqn. 29, the ratio of the RNA loss change to the original unlearned model loss change is:
@@
[…]
@@
[…]
@@
$\mathbb{P}\left[\frac{\Delta\mathcal{J}^{\text{rna}}}{\Delta\mathcal{J}^{u}}\leq 0\right]\approx\mathbb{P}\left[\frac{(\bm{g}^{\text{per}})^{\top}\bm{\delta}_{1}-\bm{g}^{\top}\bm{\delta}_{2}}{\bm{g}^{\top}\bm{\epsilon}}\leq-1\right]$ $=F(x=-1;x_{0},\gamma)$ (32)
[…]
Appendix C Empirical Validation
C.1 Empirical Validation of Section 3
In this subsection, we aim to show that the PO forgetting process (minimizing the forget-loss) can be interpreted as injecting random noise into the latent representations of forget-samples during fine-tuning.
@@ C.1 Empirical Validation of Section 3
We formalize the forgetting through the lens of noise sensitivity (Arora et al., 2018). Let $\mathbf{z}^{f}\in\mathbb{R}^{d_{l}}$ be the hidden states vector of forget-sample $\mathbf{x}^{f}$ at layer $l$ in the model $f$, where $d_{l}$ is the dimension of layer $l$. Let $g$ be the $(l+1)$-th transformer layer in model $f^{u}$. Consider a random perturbation $\bm{\xi}$ drawn from a Normal distribution $\mathcal{N}(\bm{0},\bm{I})$. The noise sensitivity of $g$ with respect to $\mathcal{N}(\bm{0},\bm{I})$ on forget-set $\mathcal{D}_{f}$, is defined as:
@@
For all unlearned models, we perform grid search for $g$ from the first to the last layer in the model. We use the WMDP-Biology forget-set to compute the noise sensitivity of layers by Eqn. 35. The max sequence length of each forget-sample is set to $512$.
@@ Results.
Discussion. We employ the noise sensitivity to validate the analysis in Section 3. However, we believe that this definition has broader potential applications. One could explore the noise sensitivity as a metric for measuring unlearning difficulty. This definition generalizes two perspectives: model difficulty and data difficulty. From the model perspective, noise sensitivity can help characterize the unlearning difficulty of specific components—such as an intermediate layer (as described in Eqn. 35), a group of layers, an entire model (e.g., Llama vs. Mistral), or more fine-grained modules in the layer such as MLP, attention patterns, or individual neurons. From the data perspective, the noise sensitivity can be used to evaluate unlearning difficulty at the level of individual samples, sub-classes, or data subsets. We leave these promising directions for future work.
The convexity assumption.
Our derivation in Section 3 is based on the assumption that the loss is locally convex w.r.t. $\mathbf{z}_{\theta}^{f}$. This assumption ensures that the Hessian matrix $\nabla^{2}_{\mathbf{z}_{\bm{\theta}}^{f}}\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f})$ is positive definite, which in turn guarantees that its trace is positive. However, if $\mathbf{z}_{\theta}^{f}$ is located at a local maximum, the Hessian would be negative definite and the sign of Eqn. 9 would flip, that is, adding noise would, in such cases, decrease the expected loss. Despite local convexity being difficult to guarantee due to the highly non-linear property of deep networks, we note that our assumption is reasonable rather than overly restrictive.
Figure 8: Distribution of loss changes for forget-samples in the WMDP-Biology and WMDP-Cyber QA datasets. The loss change is computed as the difference between the losses in the RM model and the base model.
We conduct the following empirical experiment to understand how the RM methods affect the loss of forget-samples. Specifically, we compute the loss change relative to the base model and RM models for all forget-samples in the WMDP-Biology and WMDP-Cyber QA datasets. The distribution of loss changes is shown in Figure 8. We observe that the loss changes are positive, suggesting that, in general, RM methods increase the loss of forget-samples compared to the base model. This behavior aligns with the assumption and further supports the analysis in Section 3, that adding noise typically leads to a higher loss.
C.2 Empirical Validation of Assumption 4.1
Figure 9: Distributions of changes in maximum activations (MaxAct) of generated tokens in unlearned models when conditioned on perturbed and original MMLU retain-queries.
Assumption 4.1 implies that the latent representations of generated tokens in unlearned models, given the perturbed retain-query, can be modeled as the original representation plus Gaussian noise. While this is intuitively plausible, its validity in complex LLMs might require further validation. We first discuss that Gaussian noise is a common and well-established choice for random perturbations. The Gaussian noise allows us to formally prove the core intuition that the presence of forget-tokens in retain-queries introduces noise-like perturbations to the model’s outputs.
To further support the assumption, we design the following experiment: We ask the unlearned model to generate $k=10$ tokens given the perturbed retain-query and original retain-query using greedy decoding, for all MMLU retain-queries. Perturbed MMLU retain-queries are constructed as shown in Figure 6. We then compute the difference in maximum activation (MaxAct) across latent dimensions at layer $l=7$ of each generated token, conditioned on perturbed and original MMLU retain-queries. As shown in Figure 9, the differences in MaxActs exhibit a Gaussian-like distribution. These results provide supporting empirical evidence for Assumption 4.1.